Calculate the velocity of sound through a string of length 2 m weighing 6 g 

kept stretched by a load of 600 kg.


1
Expert's answer
2021-08-10T11:48:04-0400
"v=\sqrt{\frac{T}{\mu}}\\v=\sqrt{\frac{600(9.8)(2)}{0.006}}=1400\frac{m}{s}\\"